OpenStudy (tanvirzawad):

x[n]=x[n+r] Now, if x[n]=A*cos[w*n+B] is Discrete Time periodic signal, then prove that w*r=2*pi*q, where q is integer.

OpenStudy (anonymous):

Here's a clue: the solution hinges on the fact that: \[\cos \theta=\cos(\theta \pm 2k \pi), k \in \mathbb{N}\]

OpenStudy (tanvirzawad):

thank you, CheesePi! May Allah Ta'ala guide you and us.